10. Domestic cats (Pelis cattus) are well-known junkies of the plant, Nepeta cataria, or catnip. Nevertheless, cats vary in their response to the plant, and some dont respond at all. It turns out that other cat species, including leopards (Panthera pardis), respond to catnip like house cats do some are fools for the stuff and others ignore it. This response is inherited as a relatively simple dominant Mendelian trait. In one study (this is true), researchers found that 14 of 18 leopards responded to catnip (had the dominant trait). Given these data, what is the best estimate of the frequency of wild leopards who are heterozygous for the catnip response, assuming the trait is in Hardy-Weinberg equilibrium?
Answer #1

As per Hardy-Weinberg Equilibrium, let
p = Frequency of Dominant Allele
q = Frequency of Recessive Allele
Also p + q = 1
Now,
p2 = Frequency of Homozygous Dominant Genotype
2pq = Frequency of Heterozygous Dominant Genotype
q2 = Frequency of Homozygous Recessive Genotype = 4/18 = 2/9 = 0.22

Therefore q = (2/9)1/2 = 0.47
Therefore p = 1-q = 1 - 0.47 = 0.53
Therefore Frequency of Heterozygous Dominant Genotype = 2pq = 2 * 0.53 * 0.47 = 0.4982
i.e. very close to 0.5, hence 9 out of 18 leopards can be assumed to be heterozygous for the catnip response.